give y-intercept and slope equation (-2,6) and (-9,1)
Answer by mananth(16946)

x1 y1 x2 y2
-2 6 -9 1

slope m =(y2-y1)/(x2-x1) 0.71
(1-6)/( -9-2)
(-5/-7)
m=0.71

Plug value of the slope and point (-2 ,6) in
Y =mx+b
6.00=-1.43+b
b=6-1.43
b=7.43
So the equation will be
Y =0.71 x+7.43
y intercept = 7.43